add. 3 3/4 + 3 5/6 (/ means fraction) i think it could be 7 3/10

Answer by dabanfield(803)

3 3/4 + 3 5/6
First change the mixed fractions to improper fractions:
3 3/4 = 15/4
3 5/6 = 23/6
To add these two improper fractions we need a common denominator. The smallest common denominator is 12 so we need to convert each from fourths and sixths to twelths:
15/4 = (15*3)/(4*3) = 45/12
26/6 = (26*2)/(6*2) = 52/12
Now we can add these:
45/12 + 52/12 = 97/12 = 8 1/12